# Intergalactic magnetic field lower limits up to the redshift $z\approx 3$

### v1

@@ Abstract
Large-scale intergalactic magnetic fields may contain a mixture of galactic and cosmogenic contributions, that can be probed via observations of $\gamma$-ray “echo” from sources at cosmological distances. While these contributions may be disentangled based on the difference in their redshift evolution, thus far indications of non-negligible magnetic field have been found only at low redshifts. This work aims at extending the intergalactic magnetic field constraints to redshifts $z\gtrsim 1$ using 17-year long all-sky observations of high-redshift active galactic nuclei with Fermi/LAT $\gamma$-ray telescope. Combing the Fermi/LAT measurements in the 0.1 GeV – 1 TeV energy range with the Monte Carlo simulations of the $\gamma$-ray “echo”, it is shown that the zero field strength hypothesis at $z=[0.5;3]$ is disfavoured at the $\approx 9\sigma$ significance level, yielding the lower limit of $B\gtrsim 1\times 10^{-18}$ cG for the magnetic field correlation length above 1 Mpc. It is further shown that the same data put a lower limit on the volume-filling fraction of this field of $f\gtrsim 90\%$ in the redshift range considered. It is also demonstrated that the derived limits are not substantially affected by either source flux variability or the assumed $\gamma$-ray emission attenuation model. Implications of these limits for intergalactic magnetic field origin are discussed.
@@ 1 Introduction
Most of the known galaxies and galaxy clusters are believed to contain magnetic fields generated at the early stages of the their evolution or inherited from the intergalactic space during the proto-galaxy formation [1, 2, 3]. These initial “seed” fields could have been amplified by various dynamo effects during the galaxy lifetime and further ejected into the intergalactic space along with the supernovae- or active galactic nuclei (AGN) driven outflows at redshifts $z\sim 1-4$ [4, 5, 6, 7]. While the exact volume filling factor of such outflows is uncertain, it is plausible that intergalactic space may presently contain a mixture of cosmogenic and galactic magnetic fields.
First evidences for the presence of weak intergalactic magnetic fields (IGMF) were found in $\gamma$-ray observations of distant active galactic nuclei [8]. On its way to Earth, the very-high-energy (VHE; $\gtrsim 100$ GeV) emission from these objects is subject to partial absorption in interactions with the extragalactic microwave, infrared and optical photon fields constituting the extragalactic background light (EBL) [9, 10, 11]. The power absorbed is transferred to the electromagnetic cascades, initiated in the process, and eventually re-emitted as a lower-energy gamma-ray “echo”, whose properties are sensitive to IGMF it develops within [12, 13, 14, 15]. Lack of this “echo” signal in the observed AGN and gamma-ray burst (GRB) emission has been used to derive the lower limit on the intervening IGMF [8, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Coupled with the upper limits from cosmic microwave background fluctuations and AGN radio emission Faradey rotation [15, 25, 26, 27, and references therein], these constrain IGMF strength to the $B\sim 10^{-17}-10^{-11}$ G range (for field correlation length scale in excess of a megaparsec), but so far have been insufficient to identify its nature as both cosmogenic and galactic IGMF remain consistent with the data.
Even if the present-day IGMF constrains are inconclusive, the field nature may be probed via measurements at redshifts above $z>1$, owing to the distinctly different redshift evolution of the cosmogenic and galactic fields [25, 7, 28]. However, the increasing attenuation of the intrinsic VHE emission with redshift, at $z\simeq 1$ reaching optical depth of unity already at 100 GeV energy [29], makes these challenging from the observational point of view. With this, only a handful of VHE-band AGN detections have been reported thus far at $z>0.5$ in dedicated observational campaigns with ground-based $\gamma$-ray telescopes [e.g. 30, 31, 32, 33, 34]. Furthermore, sparsity of such measurements in time results in them generally not being contemporaneous with the lower-energy “echo” constraints obtained with other instruments, dictating artificial assumptions on the AGN (non-)variability during the observational time window [e.g. 17, 18, 19, 35, 20, 21].
@@ 2.1 High-z AGN sample
The source sample used here was selected out of the known gamma-ray loud AGNs listed in the Fourth LAT AGN catalogue [4LAC; 38] with the focus on high-redshift ($z>0.5$) sources that may present detectable “echo”. First and foremost, lower limits on IGMF with $\gamma$-ray observations require an “echo” signal to be distinguishable on top of the intrinsic source emission in case of a negligibly weak magnetic field. Such a condition is more easily satisfied for hard-spectrum AGNs, where the attenuated VHE emission power may exceed that intrinsically emitted at the “echo” energies [15, 8], leading to a possibly dominant “echo” signal. With this in mind, only a subset of 4LAC AGNs with power law spectra indices $\Gamma<2$ (within the quoted uncertainties) was selected for the analysis here. To increase the chances of detecting the “echo” signal, several alternative selection criteria were used to further narrow down the list. Firstly, all sources with the confident VHE signal (4LAC detection test statistics in the VHE band $TS>25$) were selected. To include suitable source where VHE signal could have been attenuated below Fermi/LAT detection level, bright AGNs with the reported flux in the adjacent 30-100 GeV energy band $\nu F_{\nu}^{30-100}>1~\mathrm{eV/(cm^{2}~s)}$ were also added to the list. In addition to those, sources with probable VHE signal ($TS>9$) were further included. To include the sources whose VHE band detections could have been overlooked in 4LAC, VHE-detected high-redshift sources from Ref. [40] were also considered. Finally, from the sources lacking redshift estimates in 4LAC, the subset formally satisfying the above spectral index and flux requirements, for which the redshifts could have been recovered from SIMBAD astronomical database [^1] https://simbad.u-strasbg.fr/simbad/, was also included. The resulting list of sources is given in Tab.1.
Table 1: Hard-spectrum AGNs at $z>0.5$ selected here for analysis. Columns list the 4FGL catalogue name, association, redshift, spectrum type (with “PWL” standing for power law and “LP” for log-parabola correspondingly), power law spectral index, measured flux in the 30-100 GeV band, detection test statistics (TS) square root (representing the detection significance) in the VHE band as reported in 4LAC [38] and 4FGL [37] catalogues and the redshift value reference. The table is split into five sections: (1) sources with confident VHE detection ($\sqrt{TS}>5$), (2) bright sources with the flux in the adjacent 30-100 GeV energy band $\nu F_{\nu}^{30-100}>1~\mathrm{eV/(cm^{2}~s)}$, (3) sources with possible VHE signal ($\sqrt{TS}>3$), (4) VHE-detected high-redshift sources from Ref. [40] and (5) bright 4LAC sources with redshifts recovered from Refs. [42, 39, 43, 44, 45].
@@
For each of the selected sources, a time-averaged spectral energy distribution was extracted from the Fermi/LAT data over the August 2008 to July 2025 period using the Fermitools package v2.2.0 and FermiPy framework v1.2.0 [46], as described in the Fermitools [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/ and FermiPy [^3] https://fermipy.readthedocs.io/ documentation. The event selection comprised $\gamma$-ray photons corresponding to the P8R3 SOURCE type, recorded with $10^{\circ}$ from the catalogue source positions in the 0.1 GeV – 1 TeV energy range. Only events satisfying the (DATA_QUAL>0) && (LAT_CONFIG==1) good time intervals criteria were retained in the analysis. The data reduction accounted for both galactic and extragalactic diffuse emission (parametrized in the gll_iem_v07.fits and iso_P8R3_SOURCE_V2_v1.txt files provided with Fermitools) as well as the known sources from the LAT 14-year Source Catalog [4FGL-DR4; 37, 47] within $17.5^{\circ}$ from the source position with the spectral shapes taken from the catalogue. Throughout the spectrum extraction procedure, only the sources within $10^{\circ}$ from the source position were left free and only those within $5^{\circ}$ were allowed to optimize the model spectral shape in addition to its normalization. To improve the analysis sensitivity, as well as reduce the systematical uncertainties and the source confusion, the likelihood components were constructed separately for each of the Fermi/LAT PSF event types [48], restricting the usage of the worst event type quartile (evtype=4) to energies greater than 1 GeV and the second to worst (evtype=8) – to those above 0.3 GeV. All event types were used jointly in the fit, that also accounted for the energy dispersion for all components save the extragalactic diffuse one.
@@
The obtained this way source emission model, self-consistently containing the intrinsic and “echo” components, is fully parametrized by the assumed intrinsic spectrum shape and can be directly fit to the data. To this end the predicted model flux in each energy bin was compared against the measurements interpolating the “flux – likelihood” profile automatically computed by FermiPy and stored along with the extracted spectral data. Likelihood maximization itself was performed with the iminuit package [50]. The systematical uncertainties on the Fermi/LAT collection area are propagated to the fit by stretching the likelihood in each bin around its maximum by 3% below 100 GeV and $3\%+12\%\times(\log(E/\mathrm{MeV})-5)$ above this energy [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/Aeff_Systematics.html. As each energy bin is treated independently, this procedure, in principle, allows for larger systematical model-to-data deviations compared to the bracketing approach suggested for Fermi/LAT data analysis [48], conservatively reducing the analysis sensitivity to the putative “echo” signal.
@@ 3 Results
Before attempting to derive the IGMF constraints, the modelling assumptions put forward in the previous section were verified. Spatial extension of “echo” emission was estimated with CRBeam simulations and for time delays below 17 yr, corresponding to the Fermi/LAT observations duration, was found to be less than $10^{-3}$ deg in the entire IGMF strength range considered (see B). As this is well below the Fermi/LAT PSF 68% containment, the assumed $1^{\circ}$ emission cone opening and point-like treatment of the sources in Fermi/LAT analysis do not introduce any tangible uncertainty to the derived fluxes. At the same time, the constant in time “echo” flux injection assumption may not necessarily work in case of source variability within the Fermi/LAT observational time window – due to the difference in the fluence integration between the intrinsic and time-delayed “echo” emission. However, for the sources from Tab. 1 the cumulative fluence variations with respect to the mean flux do not exceed 5-10% – with exception of RX J1340.4+4410, for which “echo” prediction based on the constant flux assumption may thus introduce a sizeable systematical uncertainty (see C). Still, as is detailed below, contribution of this object to the IGMF constraint is not decisive.
Spectral energy distributions (SED) of the sources from Tab. 1, derived here, are consistent with the 4FGL records [37]; no new significant detections in the VHE range, except those already reported in 4FGL and Ref. [40], have been found. Accounting for the VHE emission attenuation, derived SEDs generally match the phenomenological power law and log-parabola shapes indicated in the catalogue. The spectral models indicated in 4LAC have been thus taken as the proxy for the intrinsic spectra [^6] In case of Ton 116, phenomenological log-parabola and exponentially cut off power law models are indistinguishable with the best-fit likelihood value difference $\Delta\log L\approx 0.06$. Though the reconstructed SED indicates no curvature below VHE energies, favouring the power law interpretation, the log-parabola model, featuring an additional curvature-related parameter, was kept here as it conservatively results in a weaker “echo” for a given IGMF strength. ; however, to conservatively restrict the total model emitted power in the VHE band for the spectra with little intrinsic curvature, an additional free exponential cut off has been introduced in all cases.
@@ 3.1 IGMF strength lower limit
The obtained best-fit likelihood profile as a function of IGMF strength is given in Fig. 1. With all sources combined, the zero IGMF hypothesis is rejected at approximately $9.1\sigma$ significance level. At 95% confidence level, the resulting sample-averaged IGMF limit at $z>0.5$ can be summarised as
@@
with $E_{\gamma}$ representing the observed “echo” emission energy, $U_{CMB}$, $T_{CMB}$ and $\epsilon_{CMB}$ denoting the cosmic microwave background energy density, temperature and photon energy, $\sigma_{T}$ being the Thomson scattering cross-section, $\sigma_{SB}$ – Stefan-Boltzmann constant and primed values corresponding to the source reference frame. As shown in the right panel of Fig. 1, the obtained $B$ limit value is comparable to that corresponding to the $z\in[1;2]$ sub-sample and 2-3 times the individual limits derived in the $z\in[1;2]$ and $z\in[2;3]$ redshift bins.
@@
However, not all the sources considered contribute to the derived limit in the same way. Several sources listed in Tab. 2 enable IGMF lower limits to be derived at the confidence level above 95% on the source-by-source basis. The strongest field limits are set by RBS 1432, Ton 116 and Ton 396, as one can also see from Fig. 1. Several sources though present the trend for the increasing best-fit likelihood value for IGMF above $B\sim 10^{-17}$ cG, suggesting a two-sided field strength measurement of $B\in[10^{-18};3\times 10^{-17}]$ cG at 95% confidence level. However, the significance of the corresponding upper limit, dominated by RBS 1432, B3 0133+388, KUV 00311-1938 and GB6 J1037+5711, does not exceed $2.8\sigma$. If these sources are excluded from the analysis, a one-sided IGMF lower limit $B\gtrsim 1\times 10^{-18}$ cG is still warranted by the rest of the sources at approximately $6\sigma$ significance level. The individual estimates of IGMF strength for the considered sources are listed in Tab. 2. For representative sources with the strongest individual contribution to the derived IGMF lower limit the spectral energy distributions and best-fit models for the minimal considered IGMF strength $B=10^{-21}$ G for are shown in Fig. 2.
@@
The choice of the VHE attenuation model, at the same time, directly affects the IGMF limits. However, the Ref. [29] model, employed here, remains a conservative choice up to $z=2$, predicting roughly 20-60% less of total VHE attenuated power at redshift $z\approx 1$ compared to the other models supported by CRBeam. Same can be said about the most recent attenuation model from Ref. [51], which predicts $7-15\%$ less attenuation depending on the redshift only at its lower uncertainty level. At redshifts $z>2$ the employed Ref. [49] model suggests attenuation similar to that of Ref. [51] with other available in CRBeam models predicting $13-80\%$ larger value. The performed above calculations may thus overestimate the “echo” power by not more than 20% – insufficient to explain the lack of the “echo” signal under the zero IGMF hypothesis. Indeed, while forcing the correspondingly lower “echo” signal in the calculations reduces the zero IGMF exclusion significance to $\approx 7.9\sigma$, it lowers the numerical limit value quoted above only twofold.

### v2

@@ Abstract
Large-scale intergalactic magnetic fields may contain a mixture of galactic and cosmogenic contributions, that can be probed via observations of $\gamma$-ray “echo” – a delayed emission from electromagnetic cascades initiated by the highest energy photons from the sources at cosmological distances. While these fields contributions may be disentangled based on the difference in their redshift evolution, thus far indications of non-negligible magnetic field have been found only at low redshifts. This work aims at extending the intergalactic magnetic field constraints to redshifts $z\gtrsim 1$ using 17-year long all-sky observations of high-redshift active galactic nuclei with Fermi/LAT $\gamma$-ray telescope. Combing the Fermi/LAT measurements in the 0.1 GeV – 1 TeV energy range with the Monte Carlo simulations of the $\gamma$-ray “echo”, it is shown that the zero field strength hypothesis at $z=[0.5;3]$ is disfavoured at the $\approx 8.6\sigma$ significance level, yielding the lower limit of $B\gtrsim 1\times 10^{-18}$ cG for the magnetic field correlation length above 1 Mpc. It is further shown that the same data put a lower limit on the volume-filling fraction of this field of $f\gtrsim 90\%$ in the redshift range considered. It is also demonstrated that the derived limits are not substantially affected by either source flux variability or the assumed $\gamma$-ray emission attenuation model. Implications of these limits for intergalactic magnetic field origin are discussed.
@@ 1 Introduction
Most of the known galaxies and galaxy clusters are believed to contain magnetic fields generated at the early stages of the their evolution or inherited from the intergalactic space during the proto-galaxy formation [41, 35, 68]. These initial “seed” fields could have been amplified by various dynamo effects during the galaxy lifetime and further ejected into the intergalactic space along with the supernovae or active galactic nuclei (AGN) driven outflows at redshifts $z\sim 1-4$ [42, 18, 46, 32]. While the exact volume filling factor of such outflows is uncertain, it is plausible that intergalactic space may presently contain a mixture of cosmogenic and galactic magnetic fields.
First evidences for the presence of weak intergalactic magnetic fields (IGMF) were found in $\gamma$-ray observations of distant active galactic nuclei [54]. On its way to Earth, the very-high-energy (VHE; $\gtrsim 100$ GeV) emission from these objects is subject to partial absorption in interactions with the cosmic microwave background as well as infrared and optical photon fields constituting the extragalactic background light (EBL) [34, 38, 60]. The power absorbed is transferred to the electromagnetic cascades, initiated in the process, and eventually re-emitted as a lower-energy gamma-ray “echo”, whose properties are sensitive to IGMF it develops within [8, 57, 51, 52]. Lack of this “echo” signal in the observed AGN and gamma-ray burst (GRB) emission has been used to derive the lower limit on the intervening IGMF [54, 62, 24, 63, 66, 4, 9, 27, 36, 67]. Coupled with the upper limits from cosmic microwave background fluctuations and AGN radio emission Faradey rotation [52, 26, 65, 58, and references therein], these constrain IGMF strength to the $B\sim 10^{-17}-10^{-11}$ G range (for field correlation length scale in excess of a megaparsec), but so far have been insufficient to identify its nature as both cosmogenic and galactic IGMF remain consistent with the data.
Even if the present-day IGMF constrains are inconclusive, the field nature may be probed via measurements at redshifts above $z>1$, owing to the distinctly different redshift evolution of the cosmogenic and galactic fields [26, 32, 64]. However, the increasing attenuation of the intrinsic VHE emission with redshift, at $z\simeq 1$ reaching optical depth of unity already at 100 GeV energy [30], makes these challenging from the observational point of view. With this, only a handful of VHE-band AGN detections have been reported thus far at $z>0.5$ in dedicated observational campaigns with ground-based $\gamma$-ray telescopes [e.g. 45, 44, 11, 3, 10]. Furthermore, sparsity of such measurements in time results in them generally not being contemporaneous with the lower-energy “echo” constraints obtained with other instruments, dictating artificial assumptions on the AGN (non-)variability during the observational time window [e.g. 24, 63, 66, 6, 4, 9].
@@ 2.1 High-z AGN sample
The source sample used here was selected out of the known gamma-ray loud AGNs listed in the Fourth LAT AGN catalogue [4LAC-DR3; 12] with the focus on high-redshift ($z>0.5$) sources that may present detectable “echo”. First and foremost, lower limits on IGMF with $\gamma$-ray observations require an “echo” signal to be distinguishable on top of the intrinsic source emission in case of a negligibly weak magnetic field. Such a condition is more easily satisfied for hard-spectrum AGNs, where the attenuated VHE emission power may exceed that intrinsically emitted at the “echo” energies [52, 54], leading to a possibly dominant “echo” signal. With this in mind, only a subset of 4LAC-DR3 AGNs with power law spectra indices111the choice of the spectral index from a power law fit here is made for the selection purposes only; a number of selected here AGNs feature curved spectral energy distributions. Still, the accepted sources with sufficient intrinsic emission curvature above few tens of GeV do not generate strong “echo” and thus do not impact the IGMF search performed here. $\Gamma<2$ (within the quoted uncertainties) was selected for the analysis here. To increase the chances of detecting the “echo” signal, several alternative selection criteria were used to further narrow down the list. Firstly, all sources with the confident VHE signal (4LAC-DR3 detection test statistics in the VHE band $TS>25$) were selected. To include suitable sources where VHE signal could have been attenuated below Fermi/LAT detection level, bright AGNs with the reported flux in the adjacent 30-100 GeV energy band $\nu F_{\nu}^{30-100}>1~\mathrm{eV/(cm^{2}~s)}$ were also added to the list. In addition to those, sources with probable VHE signal ($TS>9$) were further included. To include the sources whose VHE band detections could have been overlooked in 4LAC-DR3, VHE-detected high-redshift sources from Ref. [53] were also considered. Finally, the subset of 4LAC-DR3 sources satisfying the above spectral index and flux requirements, where the lacking redshift information could have been recovered from SIMBAD astronomical database [^2] https://simbad.u-strasbg.fr/simbad/, was also included. The resulting list of sources is given in Tab. 1.
Table 1: Hard-spectrum AGNs at $z>0.5$ selected here for analysis. Columns list the 4FGL catalogue name, association, redshift, spectrum type (with “PWL” standing for power law and “LP” for log-parabola correspondingly), power law spectral index, measured flux in the 30-100 GeV band, detection test statistics (TS) square root (representing the detection significance) in the VHE band as reported in 4LAC-DR3 [12] and 4FGL [2] catalogues and the redshift value reference. The table is split into five sections: (1) sources with confident VHE detection ($\sqrt{TS}>5$), (2) bright sources with the flux in the adjacent 30-100 GeV energy band $\nu F_{\nu}^{30-100}>1~\mathrm{eV/(cm^{2}~s)}$, (3) sources with possible VHE signal ($\sqrt{TS}>3$), (4) VHE-detected high-redshift sources from Ref. [53] and (5) bright 4LAC-DR3 sources with spectroscopic redshifts recovered from Refs. [49, 55, 1].
@@
For each of the selected sources, a time-averaged spectral energy distribution was extracted from the Fermi/LAT data over the August 2008 to July 2025 period using the Fermitools package v2.2.0 and FermiPy framework v1.2.0 [69], as described in the Fermitools [^3] https://fermi.gsfc.nasa.gov/ssc/data/analysis/ and FermiPy [^4] https://fermipy.readthedocs.io/ documentation. The event selection comprised $\gamma$-ray photons corresponding to the P8R3 SOURCE type, recorded with $10^{\circ}$ from the catalogue source positions in the 0.1 GeV – 1 TeV energy range. Only events satisfying the (DATA_QUAL>0) && (LAT_CONFIG==1) good time intervals criteria were retained in the analysis. The data reduction accounted for both galactic and extragalactic diffuse emission (parametrized in the gll_iem_v07.fits and iso_P8R3_SOURCE_V2_v1.txt files provided with Fermitools) as well as the known sources from the LAT 14-year Source Catalog [4FGL-DR4; 2, 16] within $17.5^{\circ}$ from the source position with the spectral shapes taken from the catalogue. Throughout the spectrum extraction procedure, only the sources within $10^{\circ}$ from the source position were left free and only those within $5^{\circ}$ were allowed to optimize the model spectral shape in addition to its normalization. To improve the analysis sensitivity, as well as reduce the systematical uncertainties and the source confusion, the likelihood components were constructed separately for each of the Fermi/LAT PSF event types [5], restricting the usage of the worst event type quartile (“PSF0” corresponding to evtype=4) to energies greater than 1 GeV and the second to worst (“PSF1” corresponding to evtype=8) – to those above 0.3 GeV. All event types were used jointly in the fit, that also accounted for the energy dispersion for all components save the extragalactic diffuse one.
@@
The source emission model obtained this way self-consistently contains the intrinsic and “echo” components. The model is fully parametrized by the assumed intrinsic spectral shape and can be directly fit to the data. To this end the predicted model flux in each energy bin was compared against the measurements interpolating the “flux – likelihood” profile automatically computed by FermiPy and stored along with the extracted spectral data. Likelihood maximization itself was performed with the iminuit package [23]. The systematical uncertainties on the Fermi/LAT collection area were propagated to the fit by stretching the likelihood in each bin around its maximum by 3% below 100 GeV and $3\%+12\%\times(\log(E/\mathrm{MeV})-5)$ above this energy [^6] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/Aeff_Systematics.html. As each energy bin is treated independently, this procedure, in principle, allows for larger systematical model-to-data deviations compared to the bracketing approach suggested for Fermi/LAT data analysis [5], conservatively reducing the analysis sensitivity to the putative “echo” signal.
@@ 3 Results
Before attempting to derive the IGMF constraints, the modelling assumptions put forward in the previous section were verified. Spatial extension of “echo” emission was estimated with CRBeam simulations and for time delays below 17 yr, corresponding to the Fermi/LAT observations duration, was found to be less than $10^{-3}$ deg in the entire IGMF strength range considered (see B). As this is well below the Fermi/LAT PSF 68% containment, the assumed $1^{\circ}$ emission cone opening and point-like treatment of the sources in Fermi/LAT analysis do not introduce any tangible uncertainty to the derived fluxes. At the same time, the constant in time “echo” flux injection assumption may not necessarily work in case of source variability within the Fermi/LAT observational time window – due to the difference in the fluence integration between the intrinsic and time-delayed “echo” emission. However, for the sources from Tab. 1 the cumulative fluence variations with respect to the mean flux do not exceed 5-10% – with exception of RX J1340.4+4410, for which “echo” prediction based on the constant flux assumption may introduce a sizeable systematical uncertainty (see C). Still, as is detailed below, contribution of this object to the IGMF constraint is not decisive.
Spectral energy distributions (SED) of the sources from Tab. 1, derived here, are consistent with the 4FGL records [2]; no new significant detections in the VHE range, except those already reported in 4FGL and Ref. [53], have been found. Accounting for the VHE emission attenuation, derived SEDs generally match the phenomenological power law and log-parabola shapes indicated in the catalogue. The spectral models indicated in 4LAC-DR3 have been thus taken as the proxy for the intrinsic spectra [^7] In case of Ton 116, phenomenological log-parabola and exponentially cut off power law models are indistinguishable with the best-fit likelihood value difference $\Delta\log L\approx 0.06$. Though the reconstructed SED indicates no curvature below VHE energies, favouring the power law interpretation, the log-parabola model, featuring an additional curvature-related parameter, was kept here as it conservatively results in a weaker “echo” for a given IGMF strength. ; however, to conservatively restrict the total model emitted power in the VHE band for the spectra with little intrinsic curvature, an additional free exponential cut off has been introduced in all cases.
@@ 3.1 IGMF strength lower limit
The obtained best-fit likelihood profile as a function of IGMF strength is given in Fig. 1. With all sources combined, the zero IGMF hypothesis is rejected at approximately $8.6\sigma$ significance level. At 95% confidence level, the resulting sample-averaged IGMF limit at $z>0.5$ can be summarised as
@@
with $E_{\gamma}$ representing the observed “echo” emission energy, $U_{CMB}$, $T_{CMB}$ and $\epsilon_{CMB}$ denoting the cosmic microwave background energy density, temperature and photon energy, $\sigma_{T}$ being the Thomson scattering cross-section, $\sigma_{SB}$ – Stefan-Boltzmann constant and primed values corresponding to the source reference frame. As shown in the right panel of Fig. 1, the obtained $B$ limit value is comparable to that corresponding to the $z\in[1;2]$ sub-sample and 2-3 times the individual limits derived in the $z\in[1;2]$ and $z\in[2;3]$ redshift bins. The larger in comparison limit value in the $z\in[1;2]$ bin is set by contributions of RBS 1432 and Ton 116, both displaying little curvature in the measured spectra along with the highly significant detections in the VHE band. They thus require a stronger IGMF to suppress the “echo” signal extending to higher (relative to the other sources) energies. Without these two sources, the $z\in[1;2]$ limit becomes comparable to that obtained in other redshift bins.
@@
However, not all the sources considered contribute to the derived limit in the same way. Several sources listed in Tab. 2 enable IGMF lower limits to be derived at the confidence level above 95% on the source-by-source basis. The strongest field limits are set by RBS 1432, Ton 116 and Ton 396, as one can also see from Fig. 1. Several sources though present the trend for the increasing best-fit likelihood value for IGMF above $B\sim 10^{-17}$ cG, suggesting a two-sided field strength measurement of $B\in[10^{-18};1\times 10^{-17}]$ cG at 95% confidence level. However, the significance of the corresponding upper limit, dominated by RBS 1432, KUV 00311-1938 and GB6 J0045+2127, does not exceed $2.8\sigma$. If these sources are excluded from the analysis, a one-sided IGMF lower limit $B\gtrsim 1\times 10^{-18}$ cG is still warranted by the rest of the sources at approximately $6\sigma$ significance level. The individual estimates of IGMF strength for the considered sources are listed in Tab. 2. For representative sources with the strongest individual contribution to the derived IGMF lower limit the spectral energy distributions and best-fit models for the minimal considered IGMF strength $B=10^{-21}$ G for are shown in Fig. 2.
@@
The choice of the VHE attenuation model, at the same time, directly affects the IGMF limits. However, the Ref. [30] model, employed here, remains a conservative choice up to $z=2$, predicting roughly 20-60% less of total VHE attenuated power at redshift $z\approx 1$ compared to the other models supported by CRBeam. Same can be said about the most recent attenuation models from Refs. [59, 29], with the former predicting $7-15\%$ less attenuation depending on the redshift only at its lower uncertainty level. At redshifts $z>2$ the employed Ref. [61] model suggests attenuation similar to that of Ref. [59] with other available in CRBeam models predicting $13-80\%$ larger value. The performed above calculations may thus overestimate the “echo” power by not more than 20% – insufficient to explain the lack of the “echo” signal under the zero IGMF hypothesis. Indeed, while forcing the correspondingly lower “echo” signal in the calculations reduces the zero IGMF exclusion significance to $\approx 7.6\sigma$, it lowers the numerical limit value quoted above only twofold.